# Noether’s theorem for the conditional principle of least action

### v1

@@ 3. Nother’s theorem for the problem of conditional extremum
(35) $\tilde{I}{}^{\mu}_{A}=\frac{\partial L}{\partial(\partial_{\mu}\phi^{i})}\left(\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}\psi^{\alpha}_{A}+\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\nu}\partial_{\nu}\psi^{\alpha}_{A}\right)+\left(\frac{\partial L}{\partial\phi^{i}}-\partial_{\nu}\frac{\partial L}{\partial\partial_{\nu}\phi^{i}}\right)\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}\psi^{\alpha}_{A}-\Lambda^{\mu}_{A}\,$,
which is conserved by virtue of the conditional Lagrangian equations (32).
@@
To conclude this section, let us relate these conditionally conserved currents, when the constraints are switched off, to the canonical Noether conservation laws for the unconditional extremum problem. Let no restrictions be imposed, then equations (2) reduce to identities. The gauge symmetry of such trivial constraints is the shift of the fields by an arbitrary function $\delta_{\epsilon}\phi^{i}=\epsilon^{i}$. The complete set of gauge generators reads
(36) $\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{j}=\delta^{i}_{j}\,,\quad\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{j}=0\,.$
Clearly, any global symmetry of any action can be reproduced as a specialization of the above gauge transformation, since a shift of the field by an arbitrary function can absorb any infinitesimal specific transformation. For example, for Poincaré shift of Minkowski space by the constant vector $\omega^{\mu}$, $\delta_{\omega}\phi^{i}=\omega^{\mu}\partial_{\mu}\phi^{i}$ we have $\epsilon^{i}=\omega^{\mu}\partial_{\mu}\phi^{i}$. So, the generator of this ‘‘conditional’’ symmetry reads
(37) $\psi^{i}_{\mu}=\partial_{\mu}\phi^{i}\,.$
@@
$\delta_{\omega}L(\phi,\partial\phi)=\omega^{\mu}\partial_{\mu}L=\partial_{\mu}(\Lambda^{\mu}_{\nu}\omega^{\nu}),\quad\Lambda^{\mu}_{\nu}=\delta^{\mu}_{\nu}L\,.$
@@
(38) $T^{\mu}_{\nu}=\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\partial_{\nu}\phi^{i}\,-\,L\delta^{\mu}_{\nu}\,.$
@@ 4. Examples
4.1. Example 1. Particle in central field with conserved angular momentum.
@@
(39) $S[r(t),\varphi(t)]=\int\limits_{t_{1}}^{t_{2}}\left(\frac{m\dot{r}^{2}}{2}+\frac{mr^{2}\dot{\varphi}^{2}}{2}-U(r)\right)dt,$
@@
(40) $\frac{dM}{dt}=0,\quad M=mr^{2}\dot{\varphi}\,.$
@@
(41) $\delta_{\epsilon}r=\dot{r}\epsilon-\frac{r}{2}\dot{\epsilon},\quad\delta_{\epsilon}\phi=\dot{\phi}\epsilon\,.$
@@
(42) $\frac{\delta S}{\delta\epsilon}=\frac{d}{dt}\left(\frac{m\ddot{r}r}{2}+\frac{m{\dot{r}}^{2}}{2}+U(r)+\frac{r}{2}\frac{dU}{dr}\right)=0\,,\quad\frac{dM}{dt}=0\,.$
@@
(43) $\delta_{1}\phi=\omega_{1}\,,\,\,\delta_{1}r=0\,;\qquad\delta_{2}\phi=\dot{\phi}\,\omega_{2}\,,\,\,\,\delta_{2}r=\dot{r}\,\omega_{2}\,.$
@@
(44) $\epsilon(t)=\omega=const\,.$
@@
(45) $\epsilon_{\omega}=r^{2}\omega,\quad\omega=const\,.$
@@
(46) $\delta_{\omega}r=0,\quad\delta_{\omega}\phi=\frac{M}{m}\omega\,.$
@@
(48) $K=\frac{r^{3}}{8}\left(-m\ddot{r}-\frac{dU}{dr}+mr\dot{\phi}^{2}\right)\,.$
@@
(49) $\delta_{\epsilon}r=\dot{r}\omega,\quad\delta_{\epsilon}\phi=\dot{\phi}\omega\,.$
@@
(50) $\delta_{\omega}L=\frac{dL}{dt}\omega\,.$
@@
(51) $E=\frac{m\dot{r}^{2}}{2}+\frac{mr\ddot{r}}{2}+\frac{r}{2}\frac{dU}{dr}+U(r)\,.$
@@
(52) $E=\frac{m\dot{r}{}^{2}}{2}+U_{eff}(r)\,,\quad U_{eff}(r)=U(r)+\frac{M^{2}-8Km}{2mr^{2}}.$
@@
4.2. Example 2. Particle in central field with conserved energy.
@@
(53) $\frac{dE}{dt}=0\,,\quad E=\frac{m\dot{r}^{2}}{2}+\frac{mr^{2}\dot{\phi}^{2}}{2}+U(r)\,.$
@@
(54) $\delta_{\epsilon}r=\dot{r}\epsilon+\frac{r\dot{\phi}}{\kappa}\dot{\epsilon},\quad\delta_{\epsilon}\phi=\dot{\phi}\epsilon-\frac{\dot{r}}{r\kappa}\dot{\epsilon},$
@@
(55) $\kappa=-\frac{2\dot{r}\dot{\phi}+r\ddot{\phi}}{\dot{r}}-\frac{2r\dot{\phi}U^{\prime}(r)}{m\dot{r}^{2}+mr^{2}\dot{\phi}^{2}}.$
@@
(56) $\footnotesize{\frac{\delta S}{\delta\epsilon}=-mr\dot{r}\dot{\phi}^{2}-m\ddot{r}-\frac{dU}{dr}-mr^{2}\dot{\phi}\ddot{\phi}-\frac{d}{dt}\left[\frac{1}{\kappa}\left(mr^{2}\dot{\phi}^{3}-mr\ddot{r}\dot{\phi}-r\dot{\phi}\frac{dU}{dr}-\frac{\dot{r}}{r}\frac{dM}{dt}\right)\right]=0\,;\quad\frac{dE}{dt}=0}\,.$
@@
(57) $\frac{\delta S}{\delta\epsilon}\equiv-\frac{dE}{dt}+\frac{d}{dt}\left[2E+\frac{r\dot{\phi}}{\dot{r}\kappa}\frac{dE}{dt}\right]=0\,,\qquad\frac{dE}{dt}=0\,.$
@@
Let us consider the issue about symmetries and conserved quantities for the conditionally Lagrangian system (56). The Lagrangian and the constraint do not explicitly involve time and angle, so the time shift and rotation are obvious symmetries. In the case of the usual Lagrange equations, this would immediately lead to conservation of energy and angular momentum. For a conditionally Lagrangian system, we first have to represent time shift and rotation as a specialization of gauge symmetry (54) and identify the corresponding conditional symmetry generators. Given the generators, we can construct the conserved quantities by the recipe of the previous section.
@@
(58) $\overline{\!E}=\frac{1}{\kappa}\left(mr^{2}\dot{\phi}^{3}-mr\ddot{r}\dot{\phi}-r\dot{\phi}\frac{dU}{dr}+\frac{\dot{r}}{r}\frac{dM}{dt}\right),\quad M=mr^{2}\dot{\phi}\,.$
@@
(59) $\overline{\!E}=\frac{1}{\dot{r}\kappa}(2\dot{r}\dot{\phi}+r\ddot{\phi})(m\dot{r}^{2}+mr^{2}\dot{\phi}^{2})=-\frac{dM}{dt}\left[\frac{d}{dt}\left(\frac{M}{m\dot{r}^{2}+mr^{2}\dot{\phi}^{2}}\right)\right]^{-1}\,.$
@@
(60) $\epsilon=\frac{M}{m\dot{r}^{2}+mr^{2}\dot{\phi}^{2}}\omega,\quad\omega=const\,.$
@@
(61) $\delta_{\omega}r=0,\quad\delta_{\omega}\phi=\omega\,.$
@@
(62) $\overline{\!M}=\frac{M}{\Gamma(r)}\,,\quad\Gamma(r)=\frac{E-U(r)}{E+\frac{\overline{\!E}}{2}-U(r)}\,,\quad M=mr^{2}\dot{\phi}\,.$
This constant of motion could also be found by directly integrating the relation (59) for conservation of energy. The integration is obvious as the right hand side of relation (59) is the ratio of the total derivatives, while the left hand side is a constant. Since this quantity is conserved due to the isotropy, it should be considered as the angular momentum of the conditionally Lagrangian system (56). As we see, $\overline{\!M}$ differs from the usual angular momentum $M$ by the factor $\Gamma$ involving the energy of precession $\overline{\!E}$ which is conserved due to homogeneity of time in this system. Once $\overline{\!E}\rightarrow 0$, $\Gamma\rightarrow 1$ and $\overline{\!M}\rightarrow M$. Solutions with $\overline{\!E}=0$ exist, and they obviously reproduce the usual Lagrangian dynamics where the angular momentum $\overline{\!M}$ reduces to $M$.
@@
(63) $E=\frac{m\dot{r}^{2}}{2}+U_{eff}(r)=const\,,\qquad U_{eff}(r)=\frac{\overline{\!M}{}^{2}}{2mr^{2}}\Gamma^{2}(r)+U(r)\,,\quad\Gamma(r)=\frac{E-U(r)}{E+\frac{\overline{\!E}}{2}-U(r)}\,.$
@@
(64) $t-t_{0}=\pm\int\limits_{r_{0}}^{r}\frac{dr}{\sqrt{\frac{2}{m}\left(E-U_{eff}(r,\,E,\,\overline{\!M},\,\overline{\!E})\right)}}\,.$
@@
(65) $\phi-\phi_{0}=\pm\int\limits_{r_{0}}^{r}\,\frac{\overline{\!M}\Gamma(r)}{r^{2}}\frac{dr}{\sqrt{2m\left(E-U_{eff}(r,\,E,\,\overline{\!M},\,\overline{\!E})\right)}}\,.$
@@
4.3. Proca action with imposed condition of transversality
@@
(68) $\delta_{\epsilon}A^{\mu}=\partial_{\nu}\epsilon^{\nu\mu}\,,\qquad\epsilon^{\mu\nu}=-\epsilon^{\nu\mu}\,$,
@@
(69) $\frac{\delta S}{\delta\epsilon^{\mu\nu}}=(\Box-m^{2})F_{\mu\nu}=0\,,\quad\partial_{\mu}A^{\mu}=0\,.$
@@
(70) $\delta_{\omega}A_{\mu}=\omega^{\nu}\partial_{\nu}A_{\mu}\,,\quad\omega^{\nu}=const\,.$
@@
(71) $\epsilon_{\omega}^{\mu\nu}=\omega^{\mu}A^{\nu}-\omega^{\nu}A^{\mu}\,.$
@@
(72) $\tilde{\delta}_{\omega}A^{\mu}=\omega^{\nu}\partial_{\nu}A^{\mu}-\omega^{\mu}\partial_{\nu}A^{\nu}\,.$
@@
(73) $\tilde{\delta}_{\omega}L(A,\partial A)=\omega^{\mu}\partial_{\nu}\left(\partial^{\lambda}A^{\nu}\partial_{\mu}A_{\lambda}\right)+\omega^{\nu}\partial_{\nu}L=\omega^{\mu}\partial_{\nu}\left(\partial^{\lambda}A^{\nu}\partial_{\mu}A_{\lambda}+\delta_{\mu}^{\nu}L\right)\,.$
@@
(74) $T^{\mu}_{\nu}=\delta^{\mu}_{\nu}(A_{\lambda}\Box A^{\lambda}-\frac{1}{2}m^{2}A_{\lambda}A^{\lambda}+\frac{1}{2}\partial_{\lambda}A_{\rho}\partial^{\lambda}A^{\rho})-A^{\mu}(\Box-m^{2})A_{\nu}-\partial_{\nu}A_{\lambda}\partial^{\mu}A^{\lambda}\,.$
@@
(75) $\partial_{\mu}T^{\mu}_{\nu}=-A^{\mu}(\Box-m^{2})F_{\mu\nu}-\left((\Box-m^{2})A_{\nu}\right)\partial_{\mu}A^{\mu}=0\,.$

### v2

@@ 3 Nother’s theorem for the problem of conditional extremum
$\tilde{I}{}^{\mu}_{A}=\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\left(\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}\psi^{\alpha}_{A}+\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\nu}\partial_{\nu}\psi^{\alpha}_{A}\right)+\left(\frac{\partial L}{\partial\phi^{i}}-\partial_{\nu}\frac{\partial L}{\partial\partial_{\nu}\phi^{i}}\right)\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}\psi^{\alpha}_{A}-\Lambda^{\mu}_{A}\,$, (35)
which is conserved by virtue of the conditional Lagrangian equations (32). With higher derivatives in gauge transformations and Lagrangian, this current would be defined by relations (22),(23).
@@
To conclude this subsection, let us relate these conditionally conserved currents, when the constraints are switched off, to the canonical Noether conservation laws for the unconditional extremum problem. Let no restrictions be imposed, then equations (2) reduce to identities. The gauge symmetry of such trivial constraints is the shift of the fields by an arbitrary function $\delta_{\epsilon}\phi^{i}=\epsilon^{i}$. The complete set of gauge generators reads
$\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{j}=\delta^{i}_{j}\,,\qquad\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{j}{}^{\mu}=0\,.$ (36)
Clearly, any global symmetry of any action can be reproduced as a specialization of the above gauge transformation, since a shift of the field by an arbitrary function can absorb any infinitesimal specific transformation. For example, for Poincaré shift of Minkowski space by the constant vector $\omega^{\mu}$, $\delta_{\omega}\phi^{i}=\omega^{\mu}\partial_{\mu}\phi^{i}$ we have $\epsilon^{i}_{\omega}=\omega^{\mu}\partial_{\mu}\phi^{i}$. So, the generator of this ‘‘conditional’’ symmetry reads
$\psi^{i}_{\mu}=\partial_{\mu}\phi^{i}\,.$ (37)
@@
$\delta_{\omega}L(\phi,\partial\phi)=\omega^{\mu}\partial_{\mu}L=\partial_{\mu}(\Lambda^{\mu}_{\nu}\omega^{\nu}),\qquad\Lambda^{\mu}_{\nu}=\delta^{\mu}_{\nu}L\,.$
@@
$T^{\mu}_{\nu}=\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\partial_{\nu}\phi^{i}\,-\,L\delta^{\mu}_{\nu}\,.$ (38)
@@ 4 Examples
Example 1. Particle in central field with conserved angular momentum.
@@
$S[r(t),\varphi(t)]=\int\limits_{t_{1}}^{t_{2}}\left(\frac{m\dot{r}^{2}}{2}+\frac{mr^{2}\dot{\varphi}^{2}}{2}-U(r)\right)dt,$ (39)
@@
$\frac{dM}{dt}=0,\qquad M=mr^{2}\dot{\varphi}\,.$ (40)
@@
$\delta_{\epsilon}r=\dot{r}\epsilon-\frac{r}{2}\dot{\epsilon},\qquad\delta_{\epsilon}\phi=\dot{\phi}\epsilon\,.$ (41)
@@
$\frac{\delta S}{\delta\epsilon}=\frac{d}{dt}\left(\frac{m\ddot{r}r}{2}+\frac{m{\dot{r}}^{2}}{2}+U(r)+\frac{r}{2}\frac{dU}{dr}\right)=0\,,\qquad\frac{dM}{dt}=0\,.$ (42)
@@
$\delta_{1}\phi=\omega_{1}\,,\,\,\delta_{1}r=0\,;\hskip 18.49988pt\delta_{2}\phi=\dot{\phi}\,\omega_{2}\,,\,\,\,\delta_{2}r=\dot{r}\,\omega_{2}\,.$ (43)
@@
$\epsilon(t)=\omega=const\,.$ (44)
@@
$\epsilon_{\omega}=r^{2}\omega,\qquad\omega=const\,.$ (45)
@@
$\delta_{\omega}r=0,\qquad\delta_{\omega}\phi=\frac{M}{m}\omega\,.$ (46)
@@
$K=\frac{r^{3}}{8}\left(-m\ddot{r}-\frac{dU}{dr}+mr\dot{\phi}^{2}\right)\,.$ (48)
@@
$\delta_{\epsilon}r=\dot{r}\omega,\qquad\delta_{\epsilon}\phi=\dot{\phi}\omega\,.$ (49)
@@
$\delta_{\omega}L=\frac{dL}{dt}\omega\,.$ (50)
@@
$E=\frac{m\dot{r}^{2}}{2}+\frac{mr\ddot{r}}{2}+\frac{r}{2}\frac{dU}{dr}+U(r)\,.$ (51)
@@
$E=\frac{m\dot{r}{}^{2}}{2}+U_{eff}(r)\,,\qquad U_{eff}(r)=U(r)+\frac{M^{2}-8Km}{2mr^{2}}.$ (52)
@@
Example 2. Particle in central field with conserved energy.
@@
$\frac{dE}{dt}=0\,,\qquad E=\frac{m\dot{r}^{2}}{2}+\frac{mr^{2}\dot{\phi}^{2}}{2}+U(r)\,.$ (53)
@@
$\delta_{\epsilon}r=\dot{r}\epsilon+\frac{r\dot{\phi}}{\kappa}\dot{\epsilon},\qquad\delta_{\epsilon}\phi=\dot{\phi}\epsilon-\frac{\dot{r}}{r\kappa}\dot{\epsilon},$ (54)
@@
$\kappa=-\frac{2\dot{r}\dot{\phi}+r\ddot{\phi}}{\dot{r}}-\frac{2r\dot{\phi}U^{\prime}(r)}{m\dot{r}^{2}+mr^{2}\dot{\phi}^{2}}.$ (55)
@@
$\frac{\delta S}{\delta\epsilon}=-mr\dot{r}\dot{\phi}^{2}-m\ddot{r}-\frac{dU}{dr}-mr^{2}\dot{\phi}\ddot{\phi}-\frac{d}{dt}\left[\frac{1}{\kappa}\left(mr^{2}\dot{\phi}^{3}-mr\ddot{r}\dot{\phi}-r\dot{\phi}\frac{dU}{dr}-\frac{\dot{r}}{r}\frac{dM}{dt}\right)\right]=0\,;\qquad\frac{dE}{dt}=0$
@@
$\frac{\delta S}{\delta\epsilon}\equiv-\frac{dE}{dt}+\frac{d}{dt}\left[2E+\frac{r\dot{\phi}}{\dot{r}\kappa}\frac{dE}{dt}\right]=0\,,\hskip 18.49988pt\frac{dE}{dt}=0\,.$ (57)
@@
Let us consider the issue about symmetries and conserved quantities for the conditionally Lagrangian system (56). The Lagrangian and the constraint do not explicitly involve time and angle, so the time shift and rotation are obvious symmetries. In the case of the usual Lagrange equations, this would immediately lead to conservation of energy and angular momentum. For a conditionally Lagrangian system, we first have to represent the time translation and rotation as a specialization of gauge symmetry (54) and identify the corresponding conditional symmetry generators. Given the generators, we can construct the conserved quantities by the recipe of the previous section.
@@
$\overline{\!E}=\frac{1}{\kappa}\left(mr^{2}\dot{\phi}^{3}-mr\ddot{r}\dot{\phi}-r\dot{\phi}\frac{dU}{dr}+\frac{\dot{r}}{r}\frac{dM}{dt}\right),\qquad M=mr^{2}\dot{\phi}\,.$ (58)
@@
$\overline{\!E}=\frac{1}{\dot{r}\kappa}(2\dot{r}\dot{\phi}+r\ddot{\phi})(m\dot{r}^{2}+mr^{2}\dot{\phi}^{2})=-\frac{dM}{dt}\left[\frac{d}{dt}\left(\frac{M}{m\dot{r}^{2}+mr^{2}\dot{\phi}^{2}}\right)\right]^{-1}\,.$ (59)
@@
$\epsilon=\frac{M}{m\dot{r}^{2}+mr^{2}\dot{\phi}^{2}}\omega,\qquad\omega=const\,.$ (60)
@@
$\delta_{\omega}r=0,\qquad\delta_{\omega}\phi=\omega\,.$ (61)
@@
$\overline{\!M}=\frac{M}{\Gamma(r)}\,,\qquad\Gamma(r)=\frac{E-U(r)}{E+\frac{\overline{\!E}}{2}-U(r)}\,,\qquad M=mr^{2}\dot{\phi}\,.$ (62)
This constant of motion could also be found by directly integrating the relation (59) for conservation of energy. The integration is obvious as the right hand side of relation (59) is the ratio of the total derivatives, while the left hand side is a constant. Since this quantity is conserved due to the isotropy, it should be considered as the angular momentum of the conditionally Lagrangian system (56). As we see, $\overline{\!M}$ differs from the usual angular momentum $M$ by the factor $\Gamma$ involving the energy of precession $\overline{\!E}$ which is conserved due to homogeneity of time in this system. Once $\overline{\!E}\rightarrow 0$, we have $\Gamma\rightarrow 1$ and $\overline{\!M}\rightarrow M$. Solutions with $\overline{\!E}=0$ exist, and they obviously reproduce the usual Lagrangian dynamics where the angular momentum $\overline{\!M}$ reduces to $M$.
@@
$E=\frac{m\dot{r}^{2}}{2}+U_{eff}(r)=const\,,\hskip 18.49988ptU_{eff}(r)=\frac{\overline{\!M}{}^{2}}{2mr^{2}}\Gamma^{2}(r)+U(r)\,,\qquad\Gamma(r)=\frac{E-U(r)}{E+\frac{\overline{\!E}}{2}-U(r)}\,.$ (63)
@@
$t-t_{0}=\pm\int\limits_{r_{0}}^{r}\frac{dr}{\sqrt{\frac{2}{m}\left(E-U_{eff}(r,\,E,\,\overline{\!M},\,\overline{\!E})\right)}}\,.$ (64)
@@
$\phi-\phi_{0}=\pm\int\limits_{r_{0}}^{r}\,\frac{\overline{\!M}\Gamma(r)}{r^{2}}\frac{dr}{\sqrt{2m\left(E-U_{eff}(r,\,E,\,\overline{\!M},\,\overline{\!E})\right)}}\,.$ (65)
@@
Example 3. Proca action with imposed condition of transversality
@@
$\delta_{\epsilon}A^{\mu}=\partial_{\nu}\epsilon^{\nu\mu}\,,\hskip 18.49988pt\epsilon^{\mu\nu}=-\epsilon^{\nu\mu}\,$, (68)
@@
$\frac{\delta S}{\delta\epsilon^{\mu\nu}}=(\Box-m^{2})F_{\mu\nu}=0\,,\qquad\partial_{\mu}A^{\mu}=0\,.$ (69)
@@
$\delta_{\omega}A_{\mu}=\omega^{\nu}\partial_{\nu}A_{\mu}\,,\qquad\omega^{\nu}=const\,.$ (70)
@@
$\epsilon_{\omega}^{\mu\nu}=\omega^{\mu}A^{\nu}-\omega^{\nu}A^{\mu}\,.$ (71)
@@
$\tilde{\delta}_{\omega}A^{\mu}=\omega^{\nu}\partial_{\nu}A^{\mu}-\omega^{\mu}\partial_{\nu}A^{\nu}\,.$ (72)
@@
$\tilde{\delta}_{\omega}L(A,\partial A)=\omega^{\mu}\partial_{\nu}\left(\partial^{\lambda}A^{\nu}\partial_{\mu}A_{\lambda}\right)+\omega^{\nu}\partial_{\nu}L=\omega^{\mu}\partial_{\nu}\left(\partial^{\lambda}A^{\nu}\partial_{\mu}A_{\lambda}+\delta_{\mu}^{\nu}L\right)\,.$ (73)
@@
$T^{\mu}_{\nu}=\delta^{\mu}_{\nu}(A_{\lambda}\Box A^{\lambda}-\frac{1}{2}m^{2}A_{\lambda}A^{\lambda}+\frac{1}{2}\partial_{\lambda}A_{\rho}\partial^{\lambda}A^{\rho})-A^{\mu}(\Box-m^{2})A_{\nu}-\partial_{\nu}A_{\lambda}\partial^{\mu}A^{\lambda}\,.$ (74)
@@
$\partial_{\mu}T^{\mu}_{\nu}=-A^{\mu}(\Box-m^{2})F_{\mu\nu}-\left((\Box-m^{2})A_{\nu}\right)\partial_{\mu}A^{\mu}=0\,.$ (75)